OpenStudy (anonymous):

Solve for h q=2h-2u The solution is h=

OpenStudy (lalaly):

add 2u to both sides, then divide both sides by 2

OpenStudy (anonymous):

Solve for H q=2h-2u what is the solution

OpenStudy (anonymous):

please clarify and write the solution for me

OpenStudy (turingtest):

I'm afraid I'm going to make you try to think about it, not just give away the answer. Take your original expression q=2h-2u and add 2u to BOTH SIDES what do you have now?

OpenStudy (anonymous):

q=2h+2u-2u+2u so what is the solution for H I didnt do this right did I

OpenStudy (turingtest):

no, you added 2u twice, and just to just the right side. I asked you to just ADD 2u (only once please) to BOTH sides: q+2u=2h-2u+2u can you simplify what is above (hint: what is -2u+2u )?

OpenStudy (anonymous):

0

OpenStudy (turingtest):

yes\[-2u+2u=0\]so using that, what is\[q+2u=2h-2u+2u\]after simplification?

OpenStudy (anonymous):

2

OpenStudy (turingtest):

no that is quite wrong unfortunately :( can you show me some work so I can see where you made your mistake? I don't see how you got a 2 from anywhere...

OpenStudy (anonymous):

this is so confusing for me I am doing terrible in this course, I still dont know why we need this to become a PO

OpenStudy (anonymous):

q-(2u0(2)

OpenStudy (turingtest):

I don't know what a PO is (unless you mean police officer or public official), but you have clearly been neglecting your math work for some time now. Watch this: q=2h-2u +2u +2u ------------ q+2u=2h Above is how teachers usually show it on the blackboard. remembering that \[-2u+2u=0\]do you see how that is the same as \[q=2h-2u\]add 2u to both sides: \[q+2u=2h-(2u+2u)=2h+0=2h\]? use whatever way makes the most sense to you.

OpenStudy (anonymous):

now it make ssense too me so the solution is 2h ty so much for all ur teachings (probation officer)

OpenStudy (turingtest):

NO! we are not done, and 2h is not the answer. we were asked to solve for h! that means we need h to be by itself...

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

so far we have\[2h=q+2u\]divide both sides by 2 to isolate h. try it and tell me what you get...

OpenStudy (anonymous):

1

OpenStudy (turingtest):

I'm sorry, but you need to go back to the basics. I don;t think anything I can say will help you, so this one time I will give you the answer...\[\frac{2h}{2}=\frac{q+2u}{2}\to h=\frac{q}{2}+u\]please try to study this and see what you are missing in your understanding.

OpenStudy (anonymous):

ty